To test a 0.80 kg ski helmet, it is placed at a certain height from the ground and dropped onto a force detector placed on the ground. The pulse is 0.39 N • s for an interval of 0.050 s when it hits the force detector. What is the force recorded by the force detector?


1
Expert's answer
2021-01-13T11:38:31-0500

We can find the force recorded by the force detector from the impulse momentum change equation:


"J=F\Delta t,""F=\dfrac{J}{\Delta t}=\dfrac{0.39\ N\cdot s}{0.050\ s}=7.8\ N."

Answer:

"F=7.8\ N."